OpenStudy (anonymous):

The price of a notebook was $3.35 yesterday. Today, the price fell to $3.10 . Find the percentage decrease. Round your answer to the nearest tenth of a percent.

OpenStudy (anonymous):

im not sure but i think you take 3.35-3.10=0.25 then you put 0.25/3.35=0.07462...=0.07(2dp) then times 0.07 by 100 to get the percentage so 0.07x100=7 so i think its 7%

OpenStudy (anonymous):

ive checked it its about right coz: 3.35/100x7=0.2345 then you do 3.35-0.2345=3.1155 3.1155 is about 3.10 ish :) hope this helped